OpenStudy (anonymous):

Graph -3x^2+ 12y^2 = 84. What are the domain and range.

OpenStudy (anonymous):

Is there any value for x which does not have a corresponding y-value somewhere on the graph?

OpenStudy (anonymous):

looks like a hyperbola

OpenStudy (anonymous):

1. divide throughout by 3 2. then express in standard form \[-\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]

OpenStudy (anonymous):

 I already simplified it which is \[\sqrt{7 + 1/4x^2}\]